OpenStudy (anonymous):

sqrt(x+2)+4=x can anyone help?

OpenStudy (anonymous):

\[\sqrt{x+2}+4=x\]

OpenStudy (tkhunny):

Subtract 4 to move everything except the square root to teh other side. Then suare both sides. Caution: You WILL lose something. With the Square Root, x + 2 >= 0. You no longer can see this after squaring.

OpenStudy (anonymous):

yeah i know how to do that. but when i square does x-4 turn to (x-4)(x+4) and then we use the FOIL method?

OpenStudy (tkhunny):

Why would a SQUARE turn into something else? \(x+2 = (x-4)^{2} = (x-4)(x-4)\)

OpenStudy (anonymous):

ok i went towards x-4 route and in the end i get\[\sqrt{8}+4=-6\] and \[\sqrt{5}+4=3\]

OpenStudy (tkhunny):

That makes no sense. Please include intermediate steps so we can see where you wandered off.

OpenStudy (anonymous):

uh whatever i just checked my work and i fixed something and now im figuring out the right answer